FEL Upgrade Project Weekly Brief
January 22-26, 2001


During the first week of FEL operations for the Winter Run, we demonstrated second harmonic operation of an FEL.  Operation of FELs on the even harmonics is considered forbidden by simple theory.  See the Operations section for details.


George Neil and Fred Dylla were invited to give a presentation to the Defense Science Board's panel on high
energy lasers on Thursday (Jan. 25) in Arlington.  The talk presented a status report on the kW IR Demo, the 10
kW Upgrade project, the user program and a bief discussion of the technical issuers involved in scaling to higher
powers.  It appeared to be well received and we were asked to forward a writtenn summary of the talk to the

We forwarded a draft statement of work to AFRL concerning the proposed FY 2001 effort for adding UV capability to the IR Demo upgrade.

WBS 3 (Beam Physics):

Beam loss power limits for the FEL Upgrade were examined.  Preliminary diagnostics, corrector system and configuration was passed to WBS 9-George Biallas.  Beam Physics support was provided to user run.

WBS 4 (Injector):

The second scheme for internal cesiation is ready for review. Continued work on the gun chamber design.

WBS 6 (RF):

The PO to build 16 RF Control Module by an outside vendor has been canceled and brought inside.  Requisitions are in progress to procure all of the parts.

Zone 3 - No Progress.

Zone 4 - Operated all 8 klystrons at 11 kV.  Checking 1 klystron for possible problem.  RF Control Modules and
Arc Detectors may be available in 2 weeks for RF tests.

Injector - Procurement was given the approval from the Technical Evaluation Team to use PSC, Inc to upgrade
the Quarter's HVPS's. We are changing our specification to remove the 3 phase variac in favor of multiple taps on
the power transformer and a soft on & off controller.  This should reduce the cost and save space.

Gun HVPS - Glassman believes 2 multiplier stacks are necessary for the required power output.  I am continuing
to analyze the multiplier stack circuit.

WBS 8 (I&C):

The optical BPMs for the upstream optical table, can 2, can 6, and the End-of-line Dump are installed and operational. These are prototype configurations but fully operational.  The initial results are very promising, showing ~100 micron stability - 24 hours after setting up zero (the devices are on picomotor stages and the detector is moved to FEL beam center) this was the position indication of the properly steered-up FEL beam. A four channel surface mount board has been in EECAD since Dec. awaiting layout and manufacturing. These will replace the prototypes and be available for use in each user lab. The software for "auto-steering" of the optical transport is ready but not yet checked out, this provides for user defined deadbands to be placed on any detector which in turn would use existing picomotor controls to drive the FEL beam back to center. A second hardware approach,which is nearly complete, utilizes six individual single axis
picomotor driver and a D/A to respond in real time to drifts. This system will be tested next week.

The CCD camera supplier has agreed to sell us replacement CCD sensors at 30% of the cost of new cameras,
this will result in savings of up to $20K through the life of the upgrade. We currently purchase new cameras and
only replace the smaller of the two printed circuit boards, putting the damaged CCDs aside. These will be stripped of the radiation damaged CCDs and used for the upgrade.

The first articles are due next week for the new 3" beam position monitors and the shielded beam viewers.

All of the user lab Laser Safety Systems (LSS) have  been recertified with the new option of being able to use a
hutch. This will greatly improve future response time to future changes. The new alignment mode of 10 microsecond/60 Hz is also fully functional and certified as part of the LSS, to invoke the new mode the operator simply drops the pulse width to 10usec then selects 60Hz - the rest is automatic. If a pulse width of greater than 10usec is chosen the system will trip as well as shutting of the electron beam if one were in beam mode two.

A meeting to discuss the server computer upgrades was held and a timetable was established. The new machines
will be procured as to be on site by April 2002, this will allow for six months prior to a period of "normal"
operations. The existing machines will continue to be available for initial commisiong of systems prior to beam

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o DULY Research continues to model the small dipole with the aim of achieving quality field over the good field and are
   laying out the coil configuration.
Optical Chicane Dipole (DW)
o We continued to detailing of this magnet
Arc Dipoles (GY, GX, GQ)
o The definition of the path length corrector dipole for the GY dipole as providing full wavelength of path length
   correction at 80 MeV was approved at the FEL Upgrade Meeting (Jan. 22) subject to generation of a filled out change
o We designed 4-in-hand method of winding the coils for the 180° Dipole (DY) and went through it with AES.
o We checked the layout sketches on the 180° dipole
o We defined the shape, conductor lengths and return leg thickness of the first dipole of the Arcs, providing the
   outlet of the beam to the UV machine.
QX (3.125" Quad)
o Measurement Probe: Roughing of the coil support plate was completed but an Unacceptable bow was found in
   the material.  A clamping fixture was fabricated to flatten the plate and the clamped unit sent for annealing.
o Prices for steel are being pursued by Munk/Hedrick
o Norm Vaughan will start "cleaning up" the drawings on a part-time effort (vs. GW design)
o Will Oren, head of engineering and shops has expressed a strong interest in assembling the quads here at J-Lab
   to take advantage of man-power available during that time frame (Sept. - Nov. 2001.)  A firm commitment from him
   will be pursued.  With that scenario, we would then go out with separate contracts for coils and cores.
Sextupole with H/V Correctors (Wines)
o Robin Wines is starting to construct the model of this complex magnet
H/V Correctors (Karn)
o G. Biallas reviewed layout drawings made by J. Smith and concluded that the DH style of correctors will not fit
   over the flanges on the new 3" BPMs.  The DB and DJ styles will have to be used instead.  The DH styles can still be
   used at locations involving 3" pipe.  D. Douglas provided a count of DB, DH, and DJ correctors required for the IR.
   He will work on also providing a count for the UV line.  There are here were 10 spare DB magnets in the Test Lab.
   Once the UV count is provided by Dave, we can tally the number of new magnets required and start procurement.
Trim Quad (Hiatt)
o Some effort was spent researching the feasibility of using an asymmetric iron-pole quad. We judged that such a
   magnet  would still be beyond the limits of the trim cards and that the modeling effort would be extensive.  Attention
   has returned  to firming up the Panofsky design.

WBS 10 (Wiggler):

The coil vendor developed a problem with the potting epoxy and coil shipment to PECO has been delayed.  This
should be fixed by next week.  PECO is ready with the steel and awaiting the coils.  A visit to PECO will be  off
until the coils are shipped. A rail design for Hall Probe measurements has been started. With buss bars, plumbing,
and temperature switches completed, the wiggler was powered to 150 A (close to the max current.)  The magnet
was left energized for 7 hours for "burn in".  Magnetic measurements were made using a hand-held hall probe and
it was verified all coil turns were accounted for.  It was observed that there is a 60 Gauss discrepancy between
positive and negative field pole tip pairs.  Steve Benson will explore the best cure for this DC bias field.

WBS 11 (Optics):

The prints for the top level assembly of the mirror test stand were signed off.

We pulled the quad detector assemblies for the O-BPMs in the vault and brought them upstairs for testing, then
reinstalled them.  Initial testing has been very positive, kudos to I&C for the cable runs, and software (A. Grippo) for a lot of work to get the motion control, etc working.  This promises to be a very useful system!

We performed a test of the angular repeatability of a linear stage that is being considered for the optical cavity
assembly, and it was excellent; ~ 2 urad (5 times better than specification).

We supported operations during Hot Checkout, specifically the O-BPM setup mentioned above.


Highlight of the operations this week was achieving second harmonic lasing.  This is the first time any group has been able to do this with an optical beam.  The interesting thing about this achievement is that a perfect FEL has zero gain on-axis at the second harmonic.  Second harmonic lasing of 1.5W at 2.9 microns was achieved in this case by operating in the TM01 mode and taking advantage of finite wiggler length effects.  Although of limited practical use, these tests serve to confirm our understanding of FEL physics.

Privacy and Security Notice